OpenStudy (jjohnson1):

Tracy needs to find the length and width of the equipment storage room where she works. the room is 2 meters longer than it is wide. Its perimeter is 16 meters. What is the width in meters?

OpenStudy (cj7529):

We can use algebra here If we call the width W, and the length L. From the question we find that : L=2W The perimeter is the sum of all sides, so: 2L+2W = perimeter = 16 We can no replace all L's by 2W: 16 = 2L+2W = 4W + 2W So W = 16/6 ~ 2.67m